# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ 4.2. Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue. We use this connection to motivate our Lyapunov function. In the rest of the section, we first discuss the proof strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue.
4.2.1. $M/M/\infty$ system
@@
For an $M/M/\infty$ system with arrival rate $\lambda$ and service rate $\mu$, it is known that the system is a discrete space diffusion analogue of the Ornstein–Uhlenbeck (OU) process (Chafaï, 2006). For the latter process, it is also known that it admits the Gaussian Poincaré inequality with constant $1$, which implies unit mixing rate for the OU process and also suggests that the mixing rate for $M/M/\infty$ is $\mu$. To construct an appropriate Lyapunov function, one can look at the eigenfunction that corresponds to the correct mixing rate, which is also the smallest positive real number such that the spectral measure has a jump. For $M/M/\infty$, this eigenfunction is the first non-trivial Charlier polynomial $x-\frac{\lambda}{\mu}$ (de la Iglesia, 2021). Hence, with an appropriate centering, we choose $V(q)=\left|q-\frac{\lambda}{\mu}\right|$ as the Lyapunov function so that it captures the correct eigenvalue equation $\mathcal{L}V(q)=-\mu V(q)$ for all $q$ outside of some finite set $K$ and satisfies the non-negativity condition for Lyapunov functions in Assumption 2.
4.2.2. $M/M/n$ system in the Sub Halfin-Whitt regime
@@
For a sufficiently large $n$ such that $\lambda_{n}=n-n^{1-\alpha}\geq 3$, set $V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$. If $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, set $V(q)=|q-\lambda_{n}|\forall q\leq n$. Otherwise, set
(30) $V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,q\not\in\{\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil\}\\ |\lfloor\lambda_{n}\rfloor-\lambda_{n}-1|&\textit{if }q=\lfloor\lambda_{n}\rfloor\\ |\lceil\lambda_{n}\rceil-\lambda_{n}+1|&\textit{if }q=\lceil\lambda_{n}\rceil\end{cases}$
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta(n-\lambda_{n})}$ and $\theta>0$ is a parameter to be chosen. Then, there exists $\gamma_{n},b$ with $\gamma_{n}\rightarrow 1$, such that $\mathcal{L}V\leq-\gamma_{n}V+b1_{K}$ where $b\leq 2\lambda_{n}$ and $K=\{\lambda_{n}\}$ if $\lambda_{n}\in\mathbb{Z}_{\geq 0}$ and $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ otherwise.
The proof of this Lemma is deferred to Appendix B.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue. However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 1) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix B.3.
4.3. Regime 3: Mean Field regime
@@
Let $K=\{m,...,M\}$ be a connected subset of the state space $\mathcal{S}\subset\mathbb{Z}$ of the birth-and-death process, $\nu$ be the stationary distribution of the CTMC and $\nu_{K}(x)=\nu(x)/\sum_{x\in K}\nu(x)\,\forall x\in\mathcal{S}$. Furthermore, let $\mathcal{L}$ be the generator of the birth-and-death process and assume that $\operatorname{Var}_{\nu}(f)\leq C_{P}\langle f,-\mathcal{L}f\rangle\nu\,\forall f\in\ell_{2,\nu}$, we have
(31) $\operatorname{Var}_{\nu_{K}}(f)\leq C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}\,\forall f\in\ell_{2,\nu_{K}}$
@@
In words, the above lemma allows us to obtain local mixing results if we know the mixing behavior of another system that subsumes the dynamic inside the finite set. The $M/M/\infty$ queue indeed subsumes the dynamics inside the finite set $K\in\{0,1,\ldots,n\}$. The above lemma then allows us to establish the local Poincaré constant of $\approx 1$. We call this approach the “truncation argument” to obtain a local Poincaré inequality. We then apply our Stitching Theorem (Theorem 1) and use the fact $\lambda_{n}\ll n$ to conclude that the Poincaré constant for the $M/M/n$ queue is $\approx 1$ which completes the proof. The proof details of Proposition 4 and Lemma 4 are deferred to Appendix B.4 and B.4.2 respectively.
5. Proof Details for the Super-Halfin-Whitt regime
5.1. Proof of Lemma 2
In this subsection, we will establish the $b$-weighted Poincaré results where $b$ is the positive drift term in Lemma 1. To this end, we first need to develop a toolkit to analyze the local mixing property of the birth-and-death process in a bounded state space.
@@ 5.1. Proof of Lemma 2
(Local canonical path method) For any birth-and-death process with the generator $\mathcal{L}$, a subset $K$ of the state space $\mathbb{Z}_{\geq 0}$ and the stationary distribution $\nu$, we have the following inequality holds:
(32) $\operatorname{Var}_{\nu_{K}}(f)\leq C_{L}\sum_{k:k,k+1\in K}Q_{K}(k,k+1)(f(k+1)-f(k))^{2}$
@@
Here, $C_{L}$ can also be understood as the congestion ratio, which is lower bounded by the inverse mixing rate (Levin and Peres, 2017). When $\nu_{K}$ is roughly uniform and the transitions are roughly symmetric, the constant $C_{L}$ is proportion to the sum of the path length of the path passes through an edge $e=(k,k+1)$ that maximizes the constant. As the name suggests, we prove this lemma by analyzing the path structure as follows.
@@
We emulate the proof for the canonical path method as in (Berestycki, 2020) to handle the continuous-time Markov chain. Denote $\gamma_{x,y}$ as one of the paths connecting $x,y$, note that $\mathcal{L}(x,y)>0$ if and only if $|x-y|=1$, thus all path in $K$ is a sequence of adjacent states. For any test function $f\in\ell_{2,\nu_{K}}$ and let $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\forall x\neq y$, we have:
@@
Here, (a) follows from Cauchy-Schwarz inequality and (b) follows from the interchange of sums. Hence proved. ∎
In the next step, we show that $\nu_{K}$ is a roughly uniform distribution whose support is the set $K$. This lemma lays the groundwork for us to apply the local canonical path method.
Lemma 0.
Let $\lambda_{n}=n-n^{1-\alpha},K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$ and $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$ and $\nu_{K}(x)=0$ otherwise. For $\alpha\in(1/2,1)$ and $n\geq 65$, we have
(34) $\frac{13\sqrt{3}e^{\frac{-1}{12}}n^{\alpha-1}}{320}\leq\nu_{K}(x)\leq\frac{160e^{\frac{1}{12}}}{13\sqrt{3}}n^{\alpha-1}\,\forall x\in K.$
The proof of the above lemma is deferred to Appendix B.1.2. With these results in hand, we are now ready to prove Lemma 2.
Proof of Lemma 2.
Let $\tau$ be a measure such that $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathbb{Z}_{\geq 0}$, where $b$ is defined as in (28). Note that from Lemma 1 and Lemma 2:
(35) $\tau(q)$ $\leq\frac{480\sqrt{3}}{13}e^{\frac{1}{12}}\text{ for }q=n-1,$
(36) $\tau(q)$ $\leq\frac{320e^{2}}{13\sqrt{3}}e^{\frac{1}{12}}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$

### v4

@@ 4.2 Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue and we use this connection to motivate our Lyapunov function. Thus,, we first discuss the proof strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue in the Sub-Halfin-Whitt regime.
4.2.1 $M/M/\infty$ system
@@
For an $M/M/\infty$ system with arrival rate $\lambda$ and service rate $\mu$, it is known that the system is a discrete space diffusion analogue of the Ornstein–Uhlenbeck (OU) process [12]. For the latter process, it is also known that it admits the Gaussian Poincaré inequality with constant $1$, which implies unit mixing rate for the OU process and also suggests that the mixing rate for $M/M/\infty$ is $\mu$. To construct an appropriate Lyapunov function, one can look at the eigenfunction that corresponds to the correct mixing rate, which is also the smallest positive real number such that the spectral measure has a jump. For $M/M/\infty$, this eigenfunction is the first non-trivial Charlier polynomial $x-\frac{\lambda}{\mu}$ [19]. Hence, with an appropriate centering, we choose $V(q)=\left|q-\frac{\lambda}{\mu}\right|$ as the Lyapunov function so that it captures the correct eigenvalue equation $\mathcal{L}V(q)=-\mu V(q)$ for all $q$ outside of some finite set $K$ and satisfies the non-negativity condition for Lyapunov functions in Assumption 2.
4.2.2 $M/M/n$ system in the Sub-Halfin-Whitt regime
@@
Let $\{\lambda_{n}\}_{n\in\mathbb{Z}^{+}}$ be the sequence of arrival rates of the corresponding $M/M/n$ system such that $\lambda_{n}\in(0,n)\forall n\in\mathbb{Z}^{+}$ and let $\alpha_{n}=1-\frac{\log(n-\lambda_{n})}{\log n}$. Set $V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$. If $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, set $V(q)=|q-\lambda_{n}|\forall q\leq n$. Otherwise, set
$V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,q\not\in\{\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil\}\\ |\lfloor\lambda_{n}\rfloor-\lambda_{n}-1|&\textit{if }q=\lfloor\lambda_{n}\rfloor\\ |\lceil\lambda_{n}\rceil-\lambda_{n}+1|&\textit{if }q=\lceil\lambda_{n}\rceil\end{cases}$ (35)
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta(n-\lambda_{n})}$ and $\theta>0$ is a parameter to be chosen. Then, for $\alpha_{n}\in(0,1/2)\,\forall n\in\mathbb{Z}^{+}$ and $\lim_{n\rightarrow\infty}\alpha_{n}=\alpha\in(0,1/2)$ for some constant $\alpha$, we have there exists $\gamma_{n},b$ with $\gamma_{n}\rightarrow 1$, such that
$\mathcal{L}V\leq-\gamma_{n}V+b1_{K}$ (36)
where $b\leq 2\lambda_{n}$ and $K=\{\lambda_{n}\}$ if $\lambda_{n}\in\mathbb{Z}_{\geq 0}$ and $b\leq\lceil\lambda_{n}\rceil+2$ and $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ otherwise.
The proof of this Lemma is deferred to Appendix 9.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue. However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 2) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix 9.3.
4.3 Regime 3: Mean Field regime
@@
Let $K=\{m,...,M\}$ be a connected subset of the state space $\mathcal{S}\subset\mathbb{Z}$ of the birth and death process, $\nu$ be the stationary distribution of the CTMC and $\nu_{K}(x)=\nu(x)/\sum_{x\in K}\nu(x)\,\forall x\in\mathcal{S}$. Furthermore, let $\mathcal{L}$ be the generator of the birth and death process and assume that $\operatorname{Var}_{\nu}(f)\leq C_{P}\langle f,-\mathcal{L}f\rangle\nu\,\forall f\in\ell_{2,\nu}$, we have
$\operatorname{Var}_{\nu_{K}}(f)\leq C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}\,\forall f\in\ell_{2,\nu_{K}}$ (37)
@@
In words, the above lemma allows us to obtain local mixing results if we know the mixing behavior of another system that subsumes the dynamic inside the finite set. The $M/M/\infty$ queue indeed subsumes the dynamics inside the finite set $K\in\{0,1,\ldots,n\}$. The above lemma then allows us to establish the local Poincaré constant of $\approx 1$. We call this approach the “truncation argument” to obtain a local Poincaré inequality. We then apply Theorem 2 and use the fact $\lambda_{n}\ll n$ to conclude that the Poincaré constant for the $M/M/n$ queue is $\approx 1$ which completes the proof. The proof details of Proposition 4 and Lemma 4 are deferred to Appendix 9.4 and 9.4.2 respectively.
5 Proof details of key technical lemmas
To help readers better understand how our Lyapunov-Poincaré machinery work, we shall provide a detailed proof for the Super-Halfin-Whitt with its key technical lemmas as follows. Recall in our proof sketch in Subsection 4.1 that we need to establish Lyapunov drift lemma (namely Lemma 1) and obtain the local mixing bound (Lemma 2), which consists of the following main ingredients:
  1. 1.
    Show that the stationary distribution in $K$ is roughly uniform.
  2. 2.
    Prove the local canonical path method.
  3. 3.
    Apply the canonical path method to obtain the weighted local mixing bound.
Similarly for the Sub-Halfin-Whitt regime, one can apply a similar recipe to obtain the mixing bound for the system in that regime with an appropriate Lyapunov function.
5.1 Proof of the Lyapunov drift (Lemma 1)
Proof.
$\mathcal{L}V(q)$ $=\left(\lambda_{n}e^{\theta}+\frac{n}{e^{\theta}}-(\lambda_{n}+n)\right)V(q)=\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)V(q)=-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V(q).$ (38)
Case 1 $(1\leq n\leq 7)$: In this case, we have $K=\{0,1,...,n-1\}$. From (38), we have $\mathcal{L}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V\,\,\forall x\not\in K$. If $x\in K$ then $q\leq n-1$ and since $n\leq 7$ we have that there is only a finite number of pairs $(n,q)$. Thus, we have
Case 2 $(n>7)$: We have $K=\{\lfloor 2\lambda_{n}\rfloor-n,\ldots,n\}$. This gives $V(q)=e^{\theta(n-1-q)}$ for $q\leq n-1$. So, for q<n1q<n-1, we have
For $2\lambda_{n}-n\leq q\leq n-2$ and $\alpha>1/2$, we have $V(q)=e^{\theta(n-1-q)}\leq e^{2\theta n^{1-\alpha}}$. Furthermore, we have
$\theta\leq e^{\theta}-1=\sqrt{\frac{n}{\lambda_{n}}}-1=\frac{\frac{n-\lambda_{n}}{\lambda_{n}}}{\sqrt{\frac{n}{\lambda_{n}}}+1}\leq\frac{n^{1-\alpha}}{\sqrt{n\lambda_{n}}+\lambda_{n}}=\frac{n^{-\alpha}}{\sqrt{1-n^{-\alpha}}+1-n^{-\alpha}}\leq\frac{1}{2(n^{\alpha}-1)}.$ (39)
Thus, we have $V(q)\leq e^{\frac{n^{1-\alpha}}{n^{\alpha}-1}}$ from (39). And so for $q\in K-\{n-1\}$
$\underbrace{\left(e^{\theta}-1\right)}_{O(n^{-\alpha})}\underbrace{\left(q-(2\lambda_{n}-n)\right)}_{O(n^{1-\alpha})}\underbrace{V(q)}_{O(1)}\leq\frac{1}{2(n^{\alpha}-1)}\times 2n^{1-\alpha}\times e^{\frac{n^{1-\alpha}}{n^{\alpha}-1}}\leq\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}=b(q).$
And so, we have
$\mathcal{L}V(q)\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}.$ (40)
For q=n1q=n-1, we have
$\mathcal{L}V(n-1)$ $=(\lambda_{n}+n-1)\left(e^{\theta}-1\right)\leq 2\lambda_{n}\times\frac{1}{2(n^{\alpha}-1)}=n^{1-\alpha}$
$=-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}\underbrace{V(n-1)}_{=1}+\bigg{(}n^{1-\alpha}+\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}\bigg{)}$
$\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(n-1)+n^{1-\alpha}\bigg{(}1+\frac{1}{4(n^{\alpha}-1)}\bigg{)}$
Hence we have for $q\leq n$:
$\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}.$
5.2 Proof of Lemma 2
In this subsection, we will establish the $b$-weighted Poincaré results where $b$ is the positive drift term in Lemma 1.
5.2.1 The stationary distribution inside $K$ is roughly uniform
To get a control on the finite set behavior, we first need to study the steady state distribution inside the finite set, that is to show that $\nu_{K}$ is a roughly uniform distribution whose support is the set $K$. This lemma lays the groundwork for us to apply the local canonical path method.
@@ 5.2.1 The stationary distribution inside $K$ is roughly uniform
Let $\lambda_{n}=n-n^{1-\alpha},K=\{\lfloor 2\lambda_{n}\rfloor-n,...,n-1\}$ and $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$ and $\nu_{K}(x)=0$ otherwise. For $\alpha\in(1/2,1)$ and $n\geq 110$, we have $L_{K}(n)\times n^{\alpha-1}\leq\nu_{K}(x)\leq U_{K}(n)\times n^{\alpha-1}$ where $\lim_{n\rightarrow\infty}L_{K}(n)=\lim_{n\rightarrow\infty}U_{K}(n)=1/2$.‘
The main idea of the proof is to apply Stirling’s approximation and appropriate bounds. We defer the full statement and the full proof of the lemma to Appendix 9.1.1.
5.2.2 Local canonical path lemma
To this end, we first need to develop a toolkit to analyze the local mixing property of the birth-and-death process in a bounded state space.
Lemma 6.
(Local canonical path method) For any birth and death process with the generator $\mathcal{L}$, a subset $K$ of the state space $\mathbb{Z}_{\geq 0}$ and the stationary distribution $\nu$, we have the following inequality holds:
$\operatorname{Var}_{\nu_{K}}(f)\leq C_{L}\sum_{k:k,k+1\in K}Q_{K}(k,k+1)(f(k+1)-f(k))^{2}$ (41)
@@ 5.2.2 Local canonical path lemma
Here, $C_{L}$ can also be understood as the congestion ratio, which is lower bounded by the inverse mixing rate [40]. When $\nu_{K}$ is roughly uniform and the transitions are roughly symmetric, the constant $C_{L}$ is proportion to the sum of the path length of the path passes through an edge $e=(k,k+1)$ that maximizes the constant. As the name suggests, we prove this lemma by analyzing the path structure as follows.
@@
Proof of Lemma 6: We emulate the proof for the canonical path method as in [6] to handle the continuous-time Markov chain. Denote $\gamma_{x,y}$ as one of the paths connecting $x,y$, note that $\mathcal{L}(x,y)>0$ if and only if |xy|=1|x-y|=1, thus all path in $K$ is a sequence of adjacent states. For any test function $f\in\ell_{2,\nu_{K}}$ and let $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\forall x\neq y$, we have:
@@
Here, (a) follows from the Cauchy-Schwarz inequality and (b) follows from the interchange of sums. Hence proved. $\square$
5.2.3 Proof of Lemma 2
Finally, observe that $f\in\ell_{2,\nu_{n}}$ trivially implies $f\in\ell_{2,\nu_{K}}$ and $f\in\ell_{2,\nu_{K}}$ is equivalent to $f\in\ell_{2,\tau}$ for $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathcal{S}$, given that $b$ is sufficiently well-behaved. This essentially tells us that if a test function $f$ is valid for the global distribution then it also works for the weighted distribution and the local distribution.
Proof.
Proof of Lemma 2: Let $\tau$ be a measure such that $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathbb{Z}_{\geq 0}$, where $b$ is defined as in (31). Note that from Lemma 1 and Lemma 5 and the fact that $n\geq n_{0}$, we have:
$\tau(q)$ $\leq\left(1+\frac{1}{4(n^{\alpha}-1)}\right)\frac{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\sqrt{1-2n^{-\alpha}-n^{-1}}}{2\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}}\text{ for }q=n-1,$ (43)
$\tau(q)$ $\leq\frac{e^{\frac{n^{1-2\alpha}}{1-n^{1-\alpha}}}}{1-n^{-\alpha}}\frac{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\sqrt{1-2n^{-\alpha}-n^{-1}}}{2\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$ (44)